Problem 16-01
Inventory Management

Williams & Sons last year reported sales of $128 million, cost of goods sold (COGS) of $105 and an inventory turnover ratio of 5. The company is now adopting a new inventory system. If the new system is able to reduce the firm's inventory level and increase the firm's inventory turnover ratio to 7 while maintaining the same level of sales and COGS, how much cash will be freed up? Do not round intermediate calculations. Round your answer to the nearest dollar.
Answer #1

Here, Given information:

Sales is $ 128 million

Cost of goods sold (COGS) is $ 105 million

Existing system:

Inventory turnover ratio is 5

COGS/ Average inventory = 5

Inventory = COGS / 5

Inventory = 105/5

Inventory is $ 21 million

New system:

Inventory turnover ratio is 7 (others - constant)

COGS / Average inventory = 7

Inventory = COGS/ 7

Inventory = 105/7

Inventory is $ 15 million

Free up cash:

Difference between cash spent for inventory in both systems

= Existing system - New system.

= 21- 15

= $ 6 million.